1.3 person starts walking on circular track at speed of 10,20 & 40 mtr per sec.Where they will meet on 1km circular track?
3 person starts walking on 1km long track at speed of 10,20 & 40 mtr per sec.Where they will meet ?

Answer by stanbon(75887)

3 persons start walking on 1km long track at speed of 10,20 & 40 mtr per sec. Where they will meet ?
-------
At 10 m/sec the time for one oval is 1000/10 = 100 seconds
At 20 m/sec the time for one oval is 1000/20 = 50 seconds
At 40 m/sec the time for one oval is 1000/40 = 25 seconds
------
least common multiple is 100.
They will meet at the start/finish line:
The 10 m/sec will have gone 1 lap
The 20 m/sec will have gone 2 laps
The 40 m/sec will have done 4 laps
===========================================
Cheers,
Stan H.
============